# TCKIN: A Novel Integrated Network Model for Predicting Mortality Risk in Sepsis Patients

### v1

@@ 3.3 Ablation Study
To further validate the effectiveness and importance of the components of the TCKIN model, we conducted ablation experiments. These experiments involved removing or replacing key modules to assess their impact on the overall model performance. In the first ablation study, we substituted the GRU-D (Gated Recurrent Unit with Decay) module, which is tailored to manage irregular time series data more effectively, with a standard GRU module. This change allowed us to directly compare the performance enhancements offered by GRU-D. Results, as shown in Table 4, indicate a decline in the model’s AUROC value from 0.8807 to 0.8547 and the AUPRC value from 0.5470 to 0.4957, alongside decreases in sensitivity and specificity. This underscores the critical role of GRU-D in bolstering the accuracy and stability of the predictions.
@@
Table 4: Performance Comparison of TCKIN and Its Variants on MIMIC-III and MIMIC-IV Datasets
@@
MIMIC-III TCKIN 0.8017 0.8102 0.8776 0.0721 0.5515
TCKIN without GRU-D
0.7891 0.7681 0.8519 0.0779 0.4803
TCKIN without KAN
0.8092 0.7607 0.8587 0.0738 0.5211
MIMIC-IV TCKIN 0.7875 0.8153 0.8807 0.0696 0.5470
TCKIN without GRU-D
0.8063 0.7635 0.8547 0.0748 0.4957
TCKIN without KAN
0.8173 0.7787 0.8693 0.0716 0.5304
@@ 4 Discussion
In this study, we conducted an in-depth analysis of various data features to determine their significance in a model predicting the mortality risk of sepsis patients. Our comprehensive evaluation revealed several key features that significantly influence the model’s predictive accuracy.
@@
Figure 9: The normalized contribution of the top 10 key temporal medical features in TCKIN model.
@@
Figure 10: The normalized contribution of the top 5 key constant medical features in TCKIN model.
@@
Future research should consider including a broader array of features, such as genetic markers, imaging data, and socioeconomic factors, to provide a more comprehensive assessment of a patient’s health condition and to improve the model’s predictive capabilities. Further validation and optimization could be pursued through comparative studies with other independent datasets, ensuring the model’s consistency and reliability across various settings. Through ongoing refinement and expansion, this research is poised to offer more robust support for the early detection and management of sepsis.
@@ 5 Conclusion
The Temporal-Constant KAN Integrated Network (TCKIN) model introduced in this study has demonstrated considerable superiority in predicting the mortality risk of ICU sepsis patients. By integrating constant and temporal data along with patient diagnostic ICD codes and medical ontology CCS codes, the TCKIN model achieves greater accuracy in mortality risk prediction. Experimental results show that the TCKIN model surpasses existing models on evaluation metrics such as AUROC, AUPRC, sensitivity, and specificity, underscoring its effectiveness in managing complex medical data and enhancing predictive precision. Furthermore, these results highlight the substantial benefits of the GRU-D and KAN modules in processing irregular temporal data and capturing intricate features.
This study not only confirms the effectiveness of the TCKIN model in integrating multi-source heterogeneous data but also underscores the significance of key features in prediction accuracy. These insights provide strong support for future clinical decision-making, facilitating the early identification and timely intervention for high-risk patients, thereby potentially reducing sepsis mortality rates and improving outcomes. The findings also serve as valuable references for ongoing research. By continuing to refine and expand the model, its applicability and accuracy across various clinical environments and patient demographics can be further improved. Future studies could include additional types of feature data, such as genetic and imaging data, to provide a more comprehensive assessment of patients’ health conditions and enhance the model’s predictive capabilities.

### v2

@@ 3.3 Ablation Study
To further validate the effectiveness and importance of the components of the TCKAN model, we conducted ablation experiments. These experiments involved removing or replacing key modules to assess their impact on the overall model performance. In the first ablation study, we substituted the GRU-D (Gated Recurrent Unit with Decay) module, which is tailored to manage irregular time series data more effectively, with a standard GRU module. This change allowed us to directly compare the performance enhancements offered by GRU-D. Results, as shown in Table 4, indicate a decline in the model’s AUROC value from 0.8807 to 0.8547 and the AUPRC value from 0.5470 to 0.4957, alongside decreases in sensitivity and specificity. This underscores the critical role of GRU-D in bolstering the accuracy and stability of the predictions.
@@
Table 4: Performance Comparison of TCKAN and Its Variants on MIMIC-III and MIMIC-IV Datasets (mean ± std). All results were validated using 5-fold cross-validation.
@@
MIMIC-III TCKAN 0.8017 $\pm$ 0.0169 0.8102 $\pm$ 0.0305 0.8776 $\pm$ 0.0063 0.0721 $\pm$ 0.0024 0.5515 $\pm$ 0.0227
TCKAN without GRU-D
0.7891 $\pm$ 0.0454 0.7681 $\pm$ 0.0390 0.8519 $\pm$ 0.0070 0.0779 $\pm$ 0.0021 0.4803 $\pm$ 0.0187
TCKAN without KAN
0.8092 $\pm$ 0.0264 0.7607 $\pm$ 0.0311 0.8587 $\pm$ 0.0120 0.0738 $\pm$ 0.0019 0.5211 $\pm$ 0.0225
MIMIC-IV TCKAN 0.7875 $\pm$ 0.0298 0.8153 $\pm$ 0.0318 0.8807 $\pm$ 0.0026 0.0696 $\pm$ 0.0006 0.5470 $\pm$ 0.0088
TCKAN without GRU-D
0.8063 $\pm$ 0.0359 0.7635 $\pm$ 0.0259 0.8547 $\pm$ 0.0123 0.0748 $\pm$ 0.0013 0.4957 $\pm$ 0.0259
TCKAN without KAN
0.8173 $\pm$ 0.0202 0.7787 $\pm$ 0.0114 0.8693 $\pm$ 0.0064 0.0716 $\pm$ 0.0007 0.5304 $\pm$ 0.0081
@@ 4 Discussion
This study analyzed key data features influencing sepsis mortality prediction, revealing several critical factors that enhance predictive accuracy. Previous research has demonstrated the importance of accurate and timely detection models in various medical emergencies. Similarly, these methodologies underscore the potential of machine learning in enhancing sepsis mortality prediction by integrating multiple data types.
@@
In addition to the previously discussed temporal features, it is essential to consider the effects of sepsis on cardiovascular parameters, particularly heart rate and blood pressure. Sepsis often triggers tachycardia as a compensatory response to systemic vasodilation and hypotension. Blood pressure fluctuations during sepsis are frequently characterized by episodes of hypotension, reflecting the body’s struggle to maintain adequate perfusion. Moreover, the strain on the cardiovascular system can result in arrhythmias, which serve as critical indicators of deteriorating cardiac function and are associated with increased mortality rates in septic patients. Future research should integrate these cardiovascular metrics into temporal analyses to provide a more comprehensive understanding of sepsis progression and its impact on patient outcomes.
Figure 9: The normalized contribution of the top 10 key temporal medical features in TCKAN model.
@@
Figure 10: The normalized contribution of the top 5 key constant medical features in TCKAN model.
@@
Future research should consider including a broader array of features, such as genetic markers, imaging data, and socioeconomic factors, to provide a more comprehensive assessment of a patient’s health condition and to improve the model’s predictive capabilities. Additionally, integrating advanced imaging techniques, such as spectral CT material decomposition, can significantly reduce noise and enhance data quality, which in turn may further improve the accuracy of sepsis prediction models [37]. Further validation and optimization could be pursued through comparative studies with other independent datasets, ensuring the model’s consistency and reliability across various settings. Through ongoing refinement and expansion, this research is poised to offer more robust support for the early detection and management of sepsis.
@@ 5 Conclusion
The Temporal-Constant Kolmogorov-Arnold Network (TCKAN) model presented in this study introduces several novel contributions to the field of sepsis mortality prediction. By uniquely integrating constant and temporal data with patient diagnostic ICD codes and medical ontology CCS codes, TCKAN addresses critical limitations of existing models that typically focus on a single data modality. This innovative multi-modal approach not only enhances predictive accuracy but also offers a more comprehensive and nuanced understanding of patient risk profiles, thereby setting a new standard for predictive models in critical care settings. The model outperformed existing methods across several key metrics, including AUROC, AUPRC, sensitivity, and specificity, demonstrating its robustness in handling complex medical data. The incorporation of the GRU-D and KAN modules was particularly beneficial, enhancing the model’s ability to process irregular temporal data and capture intricate features effectively.
The TCKAN model can be deployed in hospitals and Intensive Care Units (ICUs) to monitor and predict the mortality risk of sepsis patients in real-time. By providing accurate risk assessments, it also optimizes the allocation of medical resources, including the distribution of ICU beds, medical staff, and emergency equipment. Moreover, by integrating multi-source data, the model offers more detailed analyses of patient conditions, which aids in the development of personalized treatment plans and improves both treatment outcomes and patient prognosis. Additionally, the TCKAN model can function as a core component of Clinical Decision Support Systems (CDSS), helping doctors make more informed clinical decisions in complex medical environments, thereby enhancing the overall quality of healthcare services.
However, while the TCKAN model has demonstrated impressive predictive accuracy, there remains considerable scope for future research. Future studies could focus on integrating additional types of medical data, such as genetic markers, imaging data, and socioeconomic factors, to provide a more comprehensive assessment of patient health conditions. Additionally, validating the model across diverse datasets from different medical centers will be crucial to ensure its generalizability and applicability in various clinical settings. By continuously refining and expanding the TCKAN model, its utility in early detection and management of sepsis could be further enhanced, potentially leading to better patient outcomes and reduced mortality rates
Acknowledgments
This work is supported by the Yunnan Provincial Foundation for Leaders of Disciplines in Science and Technology, China under Grant 202305AC160014.
Conflict of Interest Statement
The authors declare that there are no conflicts of interest regarding the publication of this paper.